The surface area of a sphere inscribed in a regular tetrahedron is 144pi cm^2.
1. What is the radius of the sphere?

2. What is the altitude of the tetrahedron?
1
Expert's answer
2020-12-15T02:28:30-0500

The surface area of a sphere inscribed in a regular tetrahedron is "144\pi\space cm^2"

1. What is the radius of the sphere?

Surface area of a sphere "S=4\pi r^2"

"r-" radius of sphere

"r=\sqrt{\frac{S}{4\pi}}=\sqrt{\frac{144\pi}{4\pi}}=\sqrt{36}=6."

Answer: radius of the sphere is "6\space cm."

2. What is the altitude of the tetrahedron?

For a regular tetrahedron of edge length a:

Altitude "h=\sqrt{\frac{2}{3}}a"

"a=2r\sqrt{6}=12\sqrt{6}"

"h=\sqrt{\frac{2}{3}}12\sqrt{6}=24."

Answer:altitude of the tetrahedron is "24\space cm."